(i) Let V be a Banach space and let L ⊆ V be a linear manifold. Show that L, the closure of L, is a subspace of V .


(ii) Let H be a Hilbert space. Let B be an orthonormal set in H. Show that span B is dense in H if and only if B is an orthonormal basis.


1
Expert's answer
2021-02-24T07:04:00-0500

CORRECTED SOLUTION

1. Let L ⊆ V be a linear manifold, and "\bar L" be its closure. For every "x\in \bar L" there exist a sequence "\{x_n\}_{n=1}^{\infty}\subset L" which is converges to x.

Let "a, b\in\bar L" are arbitrary vectors, "a=\lim_{n\to+\infty}a_n", "b=\lim_{n\to+\infty}b_n" with "a_n, b_n\in L" for every n. Then "a_n+b_n\in L" and "\lim_{n\to+\infty}(a_n+b_n)=a+b". Hence, "a+ b\in\bar L" .

Let "\lambda" be any scalar constant, then "\lambda a_n\in L" for every n and "\lim_{n\to+\infty}\lambda a_n=\lambda\lim_{n\to+\infty} a_n =\lambda a".

Hence, "\lambda a\in\bar L" .

Since "\bar L" is a closer of L, it is shown that "\bar L" is a closed linear subspace of V.


2. if B is an orthonormal basis of H, then every element x in H may be approximated by a finite linear combinations of vectors from B, which are form span B. Hence, span B is dense.

Conversely, let span B is dense in H. Since B is an orthonormal set, for any "x\ne y \in B |x-y|=\sqrt 2" . Since H is separable, this means that B must be countable or finite. Let "B = \{v_1, v_2, ...\}" and x be an element of H. We show that the series "x_n=\sum_{k=1}^n\langle x, v_k\rangle v_k" converges to x.

Let "\varepsilon>0" and "|x-\sum_{k=1}^n\lambda_kv_k|<\varepsilon" (it is possible to find such linear combination, since span B is dense in H). Then "\langle x-x_n,v_k\rangle=\langle x,v_k\rangle-\langle x_n,v_k\rangle=\langle x,v_k\rangle-\langle x,v_k\rangle=0", hence "\langle x-x_n,x_n-\sum_{k=1}^n\lambda_kv_k\rangle=0" and, by Pythagorean theorem, "|x-\sum_{k=1}^n\lambda_kv_k|^2=|(x-x_n)+(x_n-\sum_{k=1}^n\lambda_kv_k)|^2 = |x-x_n|^2+|x_n-\sum_{k=1}^n\lambda_kv_k|^2" , which implies "|x-x_n|^2=|x-\sum_{k=1}^n\lambda_kv_k|^2-|x_n-\sum_{k=1}^n\lambda_kv_k|^2<\varepsilon^2" and "|x-x_n|<\varepsilon".

For instance, if we take "\sum_{k=1}^n\lambda_kv_k=x_{n-1}" , then we have "|x-x_n|^2=|x-x_{n-1}|^2-|x_n-x_{n-1}|^2\leq |x-x_{n-1}|^2" which implies "|x-x_n|\leq|x-x_{n-1}|".

Therefore, the sequense "|x-x_n|" is monotonous, non-negative and starting from some n is lesser than "\varepsilon". This means "x_n" converges to x and, therefore, B is an orthonormal basis in H.